A Secondary alcohol was oxidized to a ketone using hypochlorous acid. Assign the characteristic functional group stretch in each of the spectra by dragging and dropping the label into the box next to the IR signal. Then drop the compound name and classification (ketone or alcohol) of the molecule into the box under the word "compound."
Answer #1
Concepts and reason

IR spectra of two compounds is given, that is, a reactant alcohol and a product ketone. The peaks marked in the given IR spectra needs to be assigned with the correct functional group and the correct compound needs to be identified.

IR band strength is based on the type of bond formed and the reduced mass of the bond forming species. Every bond vibration has a distinct IR frequency range. Hence, identify the compound by the IR stretching frequency value.

Fundamentals

IR, that is, infrared spectroscopy is helpful in determining the type of bond and functional group present in a molecule. Vibrations produced due to stretching or bending of a band in a molecule can be detected at a specific frequency region of IR spectra. IR spectra arises in the molecule which undergo change in dipole moment while undergoing bond stretching or bending.

IR frequency range is about 500cm16670cm1500\,{\rm{c}}{{\rm{m}}^{ - 1}} - 6670\,{\rm{c}}{{\rm{m}}^{ - 1}} , the frequency in IR is represented in the units of wave number. Consider the following relation:

υ=12πkμ\upsilon = \frac{1}{{2\pi }}\sqrt {\frac{k}{\mu }}

Here,

υ=frequencyk=bondconstant,characteristicpropertyμ=reducedmass\begin{array}{l}\\\upsilon = {\rm{frequency}}\\\\k = {\rm{bond}}\,{\rm{constant, characteristic property}}\\\\\mu = {\rm{reduced mass}}\\\end{array}

In the given IR spectra (I), the peak is marked at 3350cm1{\rm{3350}}\,{\rm{c}}{{\rm{m}}^{ - {\rm{1}}}} . It corresponds to OH{\rm{O}} - {\rm{H}} stretch in the given compound.

Hence, the compound responsible for this peak is alcohol which is present in reactant diphenylmethanol.

In the given IR spectra (I), the peak is marked at 1680cm1{\rm{1680}}\,{\rm{c}}{{\rm{m}}^{ - {\rm{1}}}} . It corresponds to C=O{\rm{C = O}} stretch in the given compound.

Hence, the compound responsible for this peak is ketone which is present in the oxidized product of the reaction, that is, benzophenone.

Ans:

Hence, in the first IR spectra the characteristic functional group stretch and the name of the compound is as follows:

Name of the compound
Functional group stretching at
3350 cm --
O-H stretch
Fanclo 3350 cmen
t
Alcohol
diphenylmethanol
in kal

Name of the compound
Functional group stretching at
1680 cm-
C=0 stretch
Ketone
Benzophenone